# Identified-and-Targeted: The First Early Evidence of the Privacy-Invasive Use of Browser Fingerprinting for Online Tracking

### v1

@@ 6.1.2 No Cookies
Table 6: The number of events and data through HTTP data. The CMP is: Onetrust and the Privacy Regulation is: CCPA. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in Opt In setting. A significant difference was noticed in the Opt Out setting when comparing settings 2 and 1, suggesting that fingerprint changes may affect the tracking.
@@
Table 7: The number of events and data through HTTP data. The CMP is: Onetrust and the Privacy Regulation is: GDPR. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in Opt In setting. A significant difference of syncing events was noticed in the Opt Out setting when comparing settings 2 and 1, suggesting that fingerprint changes may affect the tracking.
@@
6.2 Cookie Restoration
@@
domain: zol.com.cn
key: z_day
value:
1. icnmo11564%3D1(language: zh_CN)
2. ixgo20%3D1(language: en_US)
@@
Takeaway [Cookie Restoration]: 1. During the data collection process for automation, we observed that certain cookies are reinstated when fingerprints remain consistent. However, when these fingerprints are altered, the values of these cookies either change or they vanish. 2. Upon manual examination, the restoration of these cookies was determined not to be related to fingerprinting.
@@
6.3 Browser Fingerprinting Data Sharing under Privacy Regulations
Takeaway [Browser Fingerprinting in GDPR]: 1. There is no conclusive evidence to suggest that Cookiebot and Didomi participate in data sharing through browser fingerprinting for user identification in both Opt Out and Opt In settings. Websites utilizing Onetrust might not be involved in data sharing activities that use browser fingerprinting to identify users in Opt In setting. 2. It appears that websites utilizing Onetrust, Quantcast, and NAI might be involved in data sharing activities that use browser fingerprinting to identify users in Opt Out setting. Quantcast and NAI might be involved in data sharing activities that use browser fingerprinting to identify users in Opt In setting.
Takeaway [Browser Fingerprinting in CCPA]: 1. There is no conclusive evidence to suggest that Cookiebot, Didomi and Quantcast participate in data sharing through browser fingerprinting for user identification in both Opt Out and Opt In settings. Websites utilizing Onetrust, and NAI might not be involved in data sharing activities that use browser fingerprinting to identify users in Opt In setting. 2. It appears that websites utilizing Onetrust, and NAI might be involved in data sharing activities that use browser fingerprinting to identify users in Opt Out setting.
7 Discussion and Limitations
Our experiment was conducted using IP addresses from two locations in the United States, both of which are located in the United States and are not subject to privacy regulations such as GDPR [15] or CCPA [16]. In regions protected by such regulations, trackers like cookies are prohibited from tracking users once they opt out. However, our experiment has revealed that advertisers may employ browser fingerprinting to track users without providing any notification. It remains uncertain whether advertisers can continue using browser fingerprinting to track users, as there is currently no established framework for auditing advertisers in this context. It’s important to note that our experiment cannot be utilized to assess advertisers’ behavior within the constraints of privacy regulations.
Another limitation of our study is that all experiments were conducted on the Linux platform. We cannot determine whether users of Windows devices, MacOS devices, or mobile devices can still be tracked by advertisers using browser fingerprinting techniques. While some of our fake fingerprint data were obtained from Windows devices, MacOS devices, or mobile devices, which we used to emulate our experimental device browsers, it would be valuable to incorporate real Windows devices, MacOS devices, or mobile devices in the “True Fingerprints" settings to gain a more comprehensive understanding.
Compared to the work of Liu et al.[33], there are notable differences in our study in both experimental design and research objectives. While our work focuses on exploring various fingerprinting settings and assessing whether different privacy regulations can constrain fingerprinting techniques, Liu et al.[33] did not involve any specific fingerprinting configurations. Instead, their research aimed to evaluate whether CMPs, websites, or advertisers comply with users’ consent choices.
8 Related Work
Some of the previous research has concentrated on the connection between online tracking and targeting.
Wills et al.[44] delved into multiple advertising platforms, paying particular attention to Google’s and Facebook’s advertising system. Their findings revealed a variety of ad types, including contextual, behavioral, location-based, and to some extent, profile-based on Google’s advertising system. Notably, while the majority of behavioral ads aligned with user interests were visible in the Ad Preference Manager, certain interests were notably missing. The team was especially taken aback by Google’s choice to showcase non-contextual ads linked to delicate subjects, such as sexual orientation, health, and financial concerns, regardless of user activity in those areas. Such ambiguity poses challenges for users trying to understand the ads presented to them. Shifting their focus to Facebook, they were unable to definitively establish a connection between external browsing habits and the ads displayed on the site. However, they observed that engaging with Facebook’s “Like" feature on third-party content had a discernible impact on ad recommendations. Alarmingly, some ads appeared to pinpoint sensitive user interests, though the details were cut off and remained unclear.
Zeng et al.[47] was carried out focusing on the influence of individual attributes, demographics, and contextual factors on ad targeting and the determination of bid values from header biddings. Drawing from data sourced from 286 participants spread over 10 websites, they determined that the website displaying the ad, the methods used for retargeting, and distinct user behaviors were pivotal in shaping targeting results. Notably, there was only a marginal association between user demographics and targeting. The investigation also highlighted that variations in bid values were largely influenced by the specific website and individual user differences. The findings emphasize the paramount importance of contextual targeting within the digital advertising realm. For advertisers, the hosting website and a user’s previous engagements with the advertiser’s platform emerged as critical pointers for relevance. In light of these observations, it is proposed that tech developers and policymakers reflect on these aspects when weighing the implications of intrusive cross-site tracking and behavioral ad targeting for effective marketing.
Cassel et al.[21] carried out an exhaustive analysis of web tracking across mobile, desktop, and browsers with an emphasis on privacy. Their findings indicated that mobile browsers received fewer tracking and advertising requests in comparison to desktop browsers. Interestingly, consistent entities seemed to be responsible for tracking and advertising on both mobile and desktop platforms. Although browsers centered on privacy successfully limited tracking and advertising solicitations, there were discernible performance variations among them. The research alluded to a trade-off: these privacy-centric browsers, while reducing tracking and ad prompts, seemed more susceptible to fingerprinting. However, it’s worth noting that this study only identified fingerprinters and did not delve deeper into the utilization of fingerprinting in advertising and targeting.
Fouad et al.[26] also investigated cookie restoration in their research. Their methodology bears similarities to ours, involving the collection of cookie data under various experimental configurations, including the use of fake fingerprints and IP addresses. However, they conducted experiments utilizing VPNs, potentially introducing increased connection latency, which could lead to inaccuracies in cookie settings. In contrast, our study utilized real IPs to mitigate this issue. Additionally, their research lacked explicit mention of any inspection process to ensure that all restored cookies resulted from fingerprinting. In our experiments, manual inspection was conducted, revealing that factors such as the current time and website data random methods could contribute to cookie restoration resembling that achieved through browser fingerprinting.
9 Conclusion
In this study, we address a crucial gap in current research and regulatory assessment practices by investigating the utilization of browser fingerprinting in advertisements for user tracking and targeting. To this end, we introduce a comprehensive framework “FPTrace” for detecting fingerprinting usage in advertising, with a particular focus on scenarios where changes in browser fingerprinting disrupt the tracking and targeting capabilities of advertisements. FPTrace undergoes rigorous evaluation within the web environment, where websites routinely collect user browsing data for sharing with advertisers. Our investigative approach entails controlled A/B experiments, data leakage of user interests, browser fingerprint manipulation, and the exploitation of advertiser bidding behavior as an informative indicator within the advertising ecosystem to discern alterations in ad campaigns. We also delve into the potential impact of altered browser fingerprints on cookie restoration. Given that advertiser bidding strategies are influenced by their prior knowledge of users, we anticipate observable bid variations when modifying browser fingerprints. Key contributions of this research include the introduction of a framework “FPTrace” tailored to the detection of advertisement changes following browser fingerprint alterations. Implemented as an extension of OpenWPM, FPTrace enables the emulation of genuine user interactions, the collection of advertiser bid data, and the recording of HTTP data. Our findings furnish compelling evidence that browser fingerprinting plays a definitive role in advertisement tracking and targeting. Bid value datasets exhibit significant disparities in trends, mean values, median values, and maximum values following changes in browser fingerprints. Furthermore, the quantity of HTTP records, encompassing HTTP chains and syncing events, undergoes substantial reduction when browser fingerprints are altered. Additionally, we explore the connection between browser fingerprinting and cookie restoration. While our results affirm the presence of browser fingerprinting information within certain cookies, they do not establish conclusive evidence of direct browser fingerprinting involvement in the cookie restoration process.
Appendix A Appendix
A.1 Analysis of HTTP data under GDPR and CCPA
@@
Table 8: The number of events and data through HTTP data. The CMP is: Cookiebot and the Privacy Regulation is: CCPA. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in both Opt Out and Opt In settings.
@@
Table 9: The number of events and data through HTTP data. The CMP is: Cookiebot and the Privacy Regulation is: GDPR. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in both Opt Out and Opt In settings.
@@
Table 10: The number of events and data through HTTP data. The CMP is: Didomi and the Privacy Regulation is: CCPA. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in both Opt Out and Opt In settings.
@@
Table 11: The number of events and data through HTTP data. The CMP is: Didomi and the Privacy Regulation is: GDPR. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in both Opt Out and Opt In settings.
@@
Table 12: The number of events and data through HTTP data. The CMP is: Quantcast and the Privacy Regulation is: CCPA. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in both Opt Out and Opt In settings.
@@
Table 13: The number of events and data through HTTP data. The CMP is: Quantcast and the Privacy Regulation is: GDPR. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. A significant difference of syncing events was noticed in both Opt Out and Opt In settings when comparing settings 2 and 1, suggesting that fingerprint changes may affect the tracking.
@@
Table 14: The number of events and data through HTTP data. The CMP is: NAI and the Privacy Regulation is: CCPA. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in Opt In setting. A significant difference was noticed in the Opt Out setting when comparing settings 2 and 1, suggesting that fingerprint changes may affect the tracking.
@@
Table 15: The number of events and data through HTTP data. The CMP is: NAI and the Privacy Regulation is: GDPR. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. A significant difference was noticed in both Opt Out and Opt In settings when comparing settings 2 and 1, suggesting that fingerprint changes may affect the tracking.

### v2

@@ A.4. Experiment Locations
Table 6. The number of events and data through HTTP data. The CMP is: Onetrust and the Privacy Regulation is: CCPA. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in Opt In setting. A significant difference was noticed in the Opt Out setting when comparing settings 2 and 1, suggesting that fingerprint changes may affect the tracking.
@@
Table 7. The number of events and data through HTTP data. The CMP is: Onetrust and the Privacy Regulation is: GDPR. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in Opt In setting. A significant difference of syncing events was noticed in the Opt Out setting when comparing settings 2 and 1, suggesting that fingerprint changes may affect the tracking.
@@
A.5. Browser Fingerprinting Data Sharing under Privacy Regulations
Takeaway [Browser Fingerprinting in GDPR]: (1) There is no conclusive evidence to suggest that Cookiebot and Didomi participate in data sharing through browser fingerprinting for user identification in both Opt Out and Opt In settings. Websites utilizing Onetrust might not be involved in data sharing activities that use browser fingerprinting to identify users in Opt In setting. (2) It appears that websites utilizing Onetrust, Quantcast, and NAI might be involved in data sharing activities that use browser fingerprinting to identify users in Opt Out setting. Quantcast and NAI might be involved in data sharing activities that use browser fingerprinting to identify users in Opt In setting.
Takeaway [Browser Fingerprinting in CCPA]: (1) There is no conclusive evidence to suggest that Cookiebot, Didomi and Quantcast participate in data sharing through browser fingerprinting for user identification in both Opt Out and Opt In settings. Websites utilizing Onetrust, and NAI might not be involved in data sharing activities that use browser fingerprinting to identify users in Opt In setting. (2) It appears that websites utilizing Onetrust, and NAI might be involved in data sharing activities that use browser fingerprinting to identify users in Opt Out setting.
A.6. Cookie Restoration Results
@@
Takeaway [Cookie Restoration]: (1) During the data collection process for automation, we observed that certain cookies are reinstated when fingerprints remain consistent. However, when these fingerprints are altered, the values of these cookies either change or they vanish. (2) Upon manual examination, the restoration of these cookies was determined not to be related to fingerprinting.
@@
A.7. Analysis of HTTP data under GDPR and CCPA
@@
Table 8. The number of events and data through HTTP data. The CMP is: Cookiebot and the Privacy Regulation is: CCPA. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in both Opt Out and Opt In settings.
@@
Table 9. The number of events and data through HTTP data. The CMP is: Cookiebot and the Privacy Regulation is: GDPR. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in both Opt Out and Opt In settings.
@@
Table 10. The number of events and data through HTTP data. The CMP is: Didomi and the Privacy Regulation is: CCPA. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in both Opt Out and Opt In settings.
@@
Table 11. The number of events and data through HTTP data. The CMP is: Didomi and the Privacy Regulation is: GDPR. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in both Opt Out and Opt In settings.
@@
Table 12. The number of events and data through HTTP data. The CMP is: Quantcast and the Privacy Regulation is: CCPA. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in both Opt Out and Opt In settings.
@@
Table 13. The number of events and data through HTTP data. The CMP is: Quantcast and the Privacy Regulation is: GDPR. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. A significant difference of syncing events was noticed in both Opt Out and Opt In settings when comparing settings 2 and 1, suggesting that fingerprint changes may affect the tracking.
@@
Table 14. The number of events and data through HTTP data. The CMP is: NAI and the Privacy Regulation is: CCPA. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. We can analyze the data from both setting 1 and setting 2. There are minimal differences observed in Opt In setting. A significant difference was noticed in the Opt Out setting when comparing settings 2 and 1, suggesting that fingerprint changes may affect the tracking.
@@
Table 15. The number of events and data through HTTP data. The CMP is: NAI and the Privacy Regulation is: GDPR. Total HTTP Records represents the total number of HTTP records collected in this settings. Total HTTP Chains represents the number of chains we built using collected HTTP data. Syncing Events represents the number of chains containing the data sharing. All data are from settings including removing cookies. A significant difference was noticed in both Opt Out and Opt In settings when comparing settings 2 and 1, suggesting that fingerprint changes may affect the tracking.